Prof. D.J. Saikia, NCRA-TIFR, Pune University Campus, Pune 411 007.

July 5, 2001

The Registrar, University of Pune, Ganeshkhind, Pune 411 007.

Dear Sir,

This is to state that I have received a copy of the thesis of Sanjay Bhatnagar, entitled “Radio Study of Galactic Supernova Remnants and the Interstellar Medium”. The thesis is being submitted to the University of Pune, towards the partial fulfillment of the award of the Ph.D. degree.

The research work presented in the thesis was carried out at the National Centre for Radio Astrophysics (NCRA), Pune, under my guidance, during the period of 1998 to 2001.

Thanking you.

Yours Sincerely

Prof. D.J. Saikia